# How Hyper-Datafication Impacts the Sustainability Costs in Frontier AI

### v1

@@ 1. Introduction
These investments would be futile without access to another critical resource: data. Contemporary frontier AI models train on datasets containing tens of trillions of data points, often scraped at scale from the internet. For example, one of the largest publicly available tokenised datasets DCLM-Pool, contains more than 240 trillion tokens (Li et al., 2024). Figure 1 illustrates the rapid growth of datasets and data volume using the datasets available on the Hugging Face Hub [^2] https://huggingface.co/datasets., a platform for hosting AI models and datasets (see Section 3).
@@ 3.1. Metadata of Hugging Face Datasets
Data collection. We collected repository-level metadata via the Hugging Face REST API [^3] Using the endpoint https://huggingface.co/api/datasets/{id} (one request per dataset ID) and retrieving attributes via the API’s expand fields., obtained dataset size statistics separately from the Hugging Face datasets-server API [^4] See https://github.com/huggingface/dataset-viewer for documentation of the datasets-server infrastructure., and extracted contextual attributes from the Hugging Face Hub Python API [^5] Accessing the public dataset catalogue at https://huggingface.co/datasets. by querying the full public dataset catalogue. These attributes are primarily self-declared and therefore exhibit lower coverage. Appendix A.1 summarises all metadata attributes and coverage.
@@ 3.2. Questionnaire for Data Workers
Data collection. We administered an English-language questionnaire via Google Forms and distributed it via the messaging application Telegram to data workers who had voluntarily joined the Telegram channel organised by a local data workers’ collective. Data collection took place in December 2025. The questionnaire comprised ten questions covering demographics, working conditions, exposure to graphic content, and employment relationships with large technology companies. The complete questionnaire is provided in Appendix B.
@@ 4.1. Environmental Costs
Figure 2 reports estimated provider-side (Left) and user-side storage (Right) energy consumption associated with Hugging Face datasets. These estimations are based on the assumptions and calculations provided in Appendix D.2. Provider-side storage energy consumption remains modest in absolute terms but increases steadily as datasets accumulate, having reached approximately 1 GWh. User-side costs are approximately three orders of magnitude larger (>>2 TWh), driven by the repeated number of downloads across a large user base. These estimates account only for actual file downloads. Streamed access does not increase download counts [^6] https://huggingface.co/docs/hub/en/datasets-download-stats. and does not generate long-term local storage. Because Hugging Face supports streaming and on-demand access, shared infrastructure likely reduces duplication relative to alternative distribution models.
Since November 2023, Hugging Face has enabled users to specify the geographic storage region for datasets [^7] https://huggingface.co/docs/hub/storage-regions.. The available regions are the US and the European Union (EU). Our results show an intense concentration of storage in the US. Of the datasets with region information (99.8%), 552,713 are hosted in the US, compared with only 316 in the EU. This is not surprising, as the majority of Hugging Face’s infrastructure is in the US.
@@ 5. Discussions
Material Costs of Data and Resource-Awareness. The narrative of data “infiniteness” behind hyper-datafication is based on the view that data has a negligible cost, and ignores the material footprint of data. In Section 4.1, we presented concrete estimations of the energy consumption and carbon footprint of data using Hugging Face datasets. Extending these estimations to all data currently used for AI development will yield even higher material costs. Measuring the cost of creating, storing, and processing datasets using existing tools like Carbontracker [^10] https://carbontracker.info/. (Anthony et al., 2020) or CodeCarbon [^11] https://codecarbon.io/. and reporting the carbon emissions in standardised datasheets as proposed by Gebru et al. (2021) are good starting points. Additionally, attaching environmental costs at each step as meta-data across the data life-cycle, as suggested by Mersy and Krishnan (2024) for carbon-provenance-based AI, can clarify the assessment of the environmental sustainability of data.
@@ Acknowledgements.
SW and RS acknowledge funding from the Independent Research Fund Denmark (DFF) under grant agreement No. 4307-00143B. RS and JK acknowledge funding from the European Union’s Horizon Europe Research and Innovation Action programme (HORIZON-CL4-2021-HUMAN-01) under grant agreement No. 101070408, project SustainML (Application Aware, Life-Cycle Oriented Model-Hardware Co-Design Framework for Sustainable, Energy Efficient ML Systems). RS further acknowledges funding from the European Union’s Horizon Europe Research and Innovation Action programme under grant agreements No. 101070284 and No. 101189771. ED acknowledges funding from the Novo Nordisk Foundation under grant agreement No. NNF24OC0090977. SM acknowledges funding from the Danish Data Science Academy (DDSA) through a DDSA Visit Grant. The authors thank the data workers who participated in the anonymous questionnaire for their time and contributions, and members of SAINTS Lab for valuable discussions.

### v2

@@ 1. Introduction
These investments would be futile without access to another critical resource: data. Contemporary frontier AI models train on datasets containing tens of trillions of data points, often scraped at scale from the internet. For example, one of the largest publicly available tokenised datasets DCLM-Pool, contains more than 240 trillion tokens (Li et al., 2024). Figure 1 illustrates the rapid growth of datasets and data volume using the datasets available on the Hugging Face Hub [^2] https://huggingface.co/datasets, a platform for hosting AI models and datasets (see Section 3).
@@ 3.1. Metadata of Hugging Face Datasets
Data collection. We collected repository-level metadata via the Hugging Face REST API [^3] Using the endpoint https://huggingface.co/api/datasets/{id} (one request per dataset ID) and retrieving attributes via the API’s expand fields., obtained dataset size statistics separately from the Hugging Face datasets-server API [^4] See https://github.com/huggingface/dataset-viewer for documentation of the datasets-server infrastructure., and extracted contextual attributes from the Hugging Face Hub Python API [^5] Accessing the public dataset catalogue at https://huggingface.co/datasets by querying the full public dataset catalogue. These attributes are primarily self-declared and therefore exhibit lower coverage. Appendix A.1 summarises all metadata attributes and coverage.
@@ 3.2. Questionnaire for Data Workers
Data collection. We administered an English-language questionnaire via an online form and distributed it via the messaging application Telegram to data workers who had voluntarily joined the Telegram channel organised by a local data workers’ collective. Data collection took place in December 2025. The questionnaire comprised ten questions covering demographics, working conditions, exposure to graphic content, and employment relationships with large technology companies. The complete questionnaire is provided in Appendix B.
@@ 4.1. Environmental Costs
Figure 2 reports estimated provider-side (Left) and user-side storage (Right) energy consumption associated with Hugging Face datasets. These estimations are based on the assumptions and calculations provided in Appendix D.2. Provider-side storage energy consumption remains modest in absolute terms but increases steadily as datasets accumulate, having reached approximately 1 GWh. User-side costs are approximately three orders of magnitude larger (>>2 TWh), driven by the repeated number of downloads across a large user base.
Since November 2023, Hugging Face has enabled users to specify the geographic storage region for datasets [^6] https://huggingface.co/docs/hub/storage-regions. The available regions are the US and the European Union (EU). Our results show an intense concentration of storage in the US. Of the datasets with region information (99.8%), 552,713 are hosted in the US, compared with only 316 in the EU. This is not surprising, as the majority of Hugging Face’s infrastructure is in the US.
@@ 5. Discussions
Material Costs of Data and Resource-Awareness. The narrative of data “infiniteness” behind hyper-datafication is based on the view that data has a negligible cost, and ignores the material footprint of data. In Section 4.1, we presented concrete estimations of the energy consumption and carbon footprint of data using Hugging Face datasets. Extending these estimations to all data currently used for AI development will yield even higher material costs. Measuring the cost of creating, storing, and processing datasets using existing tools like Carbontracker [^9] https://carbontracker.info/ (Anthony et al., 2020) or CodeCarbon [^10] https://codecarbon.io/ and reporting the carbon emissions in standardised datasheets as proposed by Gebru et al. (2021) are good starting points. Additionally, attaching environmental costs at each step as meta-data across the data life-cycle, as suggested by Mersy and Krishnan (2024) for carbon-provenance-based AI, can clarify the assessment of the environmental sustainability of data.
@@ Acknowledgements.
SW and RS acknowledge funding from the Independent Research Fund Denmark (DFF) under grant agreement No. 4307-00143B. RS and JK acknowledge funding from the European Union’s Horizon Europe Research and Innovation Action programme (HORIZON-CL4-2021-HUMAN-01) under grant agreement No. 101070408, project SustainML (Application Aware, Life-Cycle Oriented Model-Hardware Co-Design Framework for Sustainable, Energy Efficient ML Systems). RS further acknowledges funding from the European Union’s Horizon Europe Research and Innovation Action programme under grant agreements No. 101070284 and No. 101189771. ED acknowledges funding from the Novo Nordisk Foundation under grant agreement No. NNF24OC0090977. SM acknowledges funding from the Danish Data Science Academy (DDSA) through DDSA Visit Grant No. 2025-5673. The authors thank the data workers who participated in the anonymous questionnaire for their time and contributions, and members of SAINTS Lab for valuable discussions.